One number is 14 more than another number. Their product is 351. What re the numbers?
Answer by fractalier(6550)

Let's call the first number x. The other is then x + 14.
Then we have
x(x + 14) = 351
x^2 + 14x - 351 = 0
(x + 27)(x - 13) = 0
so that
x = -27 or x = 13
thus the larger numbers are
x+14 = -13 or x+14 = 27